# Knowledge Distillation on Spatial-Temporal Graph Convolutional Network for Traffic Prediction

### v2

@@ Abstract
Traffic prediction Spatial-temporal graph knowledge distillation Spatial-temporal graph neural network pruning Model compression Teacher-student architecture
@@ 1 Introduction
Traffic prediction has gained significant importance due to the increasing use of vehicles in various contexts such as transportation of goods, city taxis, personal cars, buses, etc. [1, 2]. The data used for traffic prediction is obtained by monitoring stations in various urban areas. These stations record various parameters, including vehicle speed, traffic density, traffic flow, and …, at specific time periods. These records can be modeled as temporal graphs through preprocessing [1]. Each graph node represents the traffic data of each monitoring station and connections between these stations indicate graph edges. With this data in hand, spatio-temporal graph neural networks will be capable of predicting future traffic conditions for each of the graph nodes at desired time intervals. In real-world traffic prediction, dealing with a large number of nodes can be computationally expensive and may demand substantial hardware resources. Our objective is to propose an idea to enhance the execution time of spatio-temporal graph convolutional networks (ST-GCNs) within the domain of traffic prediction. These networks serve as robust tools in deep learning and complex data analysis, dedicated to predicting relationships among various elements within spatio-temporal systems.
Leveraging knowledge distillation [4], we formulate a network that, owing to its reduced number of parameters and subsequently lower execution time, can accurately predict traffic comparable to that of a complex network. Initially, through knowledge distillation, we introduce a cost function that trains a network with fewer parameters, denoted as the ”student,” by drawing insights from a comprehensive model known as the ”teacher.” This training occurs under the consideration of spatial-temporal structures and their temporal relationships in graphs. The objective is to equip the student network with accuracy akin to the teacher model. While knowledge distillation effectively addresses the challenge of training a compact network, it falls short in providing guidance on the architecture of the student network. To overcome this, we introduce a pruning algorithm capable of identifying and eliminating insignificant parameters in a complex teacher network. In leveraging knowledge distillation, we view parameter learnability as a solution for elimination. The algorithm adeptly prunes the network while jointly training it using knowledge distillation.
We evaluate our proposed cost function and algorithm using two real-world traffic datasets: PeMSD7, comprising 228 nodes, and PeMSD8, consisting of 170 nodes. The results demonstrate that the student trained with our cost function outperforms previous approaches in terms of knowledge distillation and execution time. Additionally, we showcase that our pruning algorithm effectively determines the architecture of the student network, enhancing learnability during training with knowledge distillation compared to when the algorithm does not employ knowledge distillation.
In the following subsection, we will provide further explanation of the proposed knowledge distillation and pruning algorithm.
1.1 Knowledge Distillation and Pruning
Knowledge distillation (KD) and Pruning are two complementary techniques employed in this study for model compression and performance enhancement. KD, serving as a model compression approach, involves transferring knowledge from a complex teacher model to a simpler student model, with the aim of reducing computational overhead [4, 8, 9, 10]. This process entails training the smaller network in an optimized manner to mitigate potential accuracy loss resulting from parameter reduction [4].
In the specific context of ST-GNNs for traffic prediction, KD offers advantages such as reducing model size and the number of parameters, with the teacher model transferring spatial and temporal complexities to the student model for accelerated learning while maintaining accuracy.
Knowledge distillation (KD) is categorized into response-based, feature-based, and relation-based methods [11, 12, 13, 14, 15, 16, 17]. Response distillation focuses on the final layer’s output, while feature distillation targets intermediate layers. This paper adopts both response-based and feature-based knowledge distillation methods. Additionally, KD can be classified based on distillation strategy into offline, online, and self-distillation [18, 19, 20]. The chosen strategy in this work is offline distillation, where both student and teacher networks are updated. An inherent challenge in KD is the selection of the student network structure, addressed in this study by intentionally reducing parameters in the student network compared to the teacher model. The paper utilizes pruning methods to achieve this reduction, and the remaining student network parameters’ learnability is a key criterion for employing KD.
Pruning, as a separate neural network compression technique, is employed to reduce model complexity and enhance performance by eliminating unnecessary parameters and connections [5, 21].
Specifically, in this study, pruning is applied to filters in hidden layers, involving the removal of weights or connections with the lowest importance based on specified thresholds derived from methods like gradient analysis or entropy information [6]. The integration of KD into the pruning process establishes criteria for assessing the impact of network parameters on model accuracy and their ability to learn from the distilled data provided by the teacher model. The evaluation of parameter competency becomes a crucial criterion in the joint application of knowledge distillation and pruning.
@@ 2.1 Traffic Prediction and Spatio-Temporal Graph Neural Networks
Fast and accurate traffic prediction is essential for urban traffic control and management, particularly for medium and long-term forecasts. Traditional methods often struggle with these predictions due to the complexities of traffic flows and the neglect of spatial and temporal dependencies. To address this, deep networks capable of handling temporal and spatial data have been proposed. In [22], the paper discusses the significance of traffic network prediction and its applications, such as network monitoring, resource management, and threat detection. Various recurrent neural network architectures, including standard RNN, LSTM, and GRU, are explored for network traffic prediction. In [23], the development and evaluation of short-term traffic prediction models using bidirectional and unidirectional long short-term memory (LSTM) neural networks are examined. Bidirectional LSTM (BiLSTM) models outperform unidirectional LSTM (Uni-LSTM) models, especially for speed and traffic flow prediction.
A combination of uni-LSTM and BiLSTM architectures also enhances 15-minute predictions. However, LSTM and RNN networks have high execution times due to numerous parameters. For low-latency traffic prediction, spatio-temporal graph neural networks are introduced. These networks utilize graph convolution layers for spatial processing and convolution layers for capturing temporal correlations, offering reduced execution time and comparable or better accuracy.
In [3], the spatio-temporal graph convolutional network (ASTGCN) model is introduced to predict traffic flow. ASTGCN addresses challenges in modeling spatial-temporal correlations in traffic data, using three components to capture recent, daily, and weekly dependencies. Spatial-temporal attention mechanisms and convolution layers effectively capture traffic data patterns, outperforming other methods in traffic flow prediction. In  [1], the spatio-temporal graph convolutional networks (STGCN) framework is presented for time series prediction in traffic management. STGCN employs a graph-based convolutional structure, which enhances training efficiency and captures comprehensive spatial-temporal correlations. It consists of two spatio-temporal convolution blocks and a fully connected layer for making predictions. This approach outperforms other methods on real traffic datasets, providing a more efficient solution for traffic forecasting.
@@ 2.2 Knowledge Distillation
Knowledge Distillation, also known as Teacher-Student Learning, is a method introduced in the paper [4] in 2015. In [4], the authors introduced knowledge distillation, a method that transfers knowledge from a complex and accurate neural network (the ”teacher”) to a smaller and more efficient neural network (the ”student”). The aim is to align the behavior and predictions of the student network with the teacher network, allowing the student to benefit from the rich knowledge of the teacher, leading to improved performance and reduced computational resources required for predictions. Initially, knowledge distillation was applied to training smaller models from complex ones, with the teacher network being a complex, accurate neural network with many parameters, and the student network being trained to align its outputs with the teacher’s, resulting in improved performance and efficiency. In further research, knowledge distillation extended to graph networks and spatiotemporal graph networks.
In [25], the focus is on knowledge transfer in Convolutional Neural Networks (CNN) and graph convolutional networks (GCN), introducing a novel approach for knowledge transfer from a GCN teacher model to a GCN student model, leading to a smaller model with improved execution time, particularly suited for dynamic graph models. In [26], the objective is to train a simpler network through knowledge distillation from multiple GCN teacher models with different tasks, extending to distill spatial graph information from hidden convolutional layers. This approach empowers the student model to excel in various tasks without additional labeled data. In [27], knowledge distillation is extended to spatio-temporal graph neural networks for modeling spatial and temporal data in human body position videos, utilizing various techniques, including minimizing loss, leveraging spatial-temporal relations, and using the gradient rejuvenation method to optimize the student model with distilled knowledge from the teacher.
@@ 2.3 Pruning and Fine-Tuning
In 1990, the concept of neural network pruning was first presented in the paper by [5]. Pruning involves eliminating unnecessary connections or neurons from trained neural networks to enhance computational efficiency without compromising network performance significantly. The authors presented a method for identifying connections safe to prune using the Hessian matrix associated with the network’s error function. This approach aims to reduce the parameters in large neural networks, lowering the risk of overfitting, decreasing computational resource requirements, and facilitating deployment on resource-constrained devices. Pruning connections has become an essential tool for compressing neural networks and optimizing models for efficient deployment across various applications.
In 2019, [6] introduced a method for estimating the importance of layers and weights in neural networks, enabling the removal of unimportant weights and reducing network size. This method accurately assesses the importance of each layer and network parameter, followed by the removal of unimportant weights. The iterative pruning and retraining stages enhance performance and reduce model complexity. Removing unimportant weights reduces network size, leading to reduced computational resource consumption and improved model execution speed. This approach offers a substantial improvement in compressing large network models.
@@ 3 Proposed Methods
Traffic prediction, involves predicting future values of key traffic conditions (e.g., speed or traffic flow) for the next $h$ time steps based on $M$ prior traffic observations. The observation vector $V_{t}\in\mathbb{R}^{N}$ at time step $t$ consists of data from $N$ monitoring stations, encompassing information like speed and traffic flow. Each time step generates a new observation vector. In this paper, we model the traffic network as graphs defined across time steps. The $N$ components of observation $V_{t}$ are interconnected through pairwise connections in the graph, making the data points $V_{t}$ graph signals in a directed graph $G$ with weights $W_{ij}$. At time step $t$, data from $N$ monitoring stations is represented as a graph $G_{t}=(V_{t},E,W)$, where $V_{t}$ contains features from $N$ nodes (e.g., road segments) with speed as the chosen criterion. The edge set $E$ depicts connections between stations, defined based on distance criteria, reflecting the influence of stations on each other. Weights $W\in\mathbb{R}^{N\times N}$ in the weighted adjacency matrix signify the relationships between stations, often determined by spatial disparities.
@@
   PeMSD7 Models Parameters Hidden Blocks Channel Test Time (s) FLOPS Teacher 296,426 [1, 32, 64][64, 32, 128] 2.556 40,636,466,453 Pruning Base Model 39,290 [1, 8, 16][16, 8, 32] 0.810 5,659,617,749 Student 7,766 [1, 2, 4][4, 2, 8] 0.441 1,003,700,933
3.1 Neural Architecture
The ST-GCN[1], designed for traffic prediction, processes spatial and temporal data. Graph convolution layers capture spatial relationships and node features over time. Convolutional layers model temporal correlations, focusing on sequential patterns and changes between graphs. The architecture comprises two blocks with temporal and spatial layers, and an output block transforms information into the final output. The teacher and student networks share an identical number of blocks and layers. Any variations in parameters result from differences in channel numbers within the temporal and spatial layers of the hidden layer. The complex ST-GCN teacher network achieves high accuracy with 333,604 and 296,426 parameters for PeMSD7 and PeMSD8 datasets ( Table 1 and Figure 1 ). Despite accuracy, its high parameter count results in significant computational cost. The challenge lies in balancing accuracy and computational cost, crucial for selecting models. To address the need for a lighter model, the student network is a scaled-down version of the teacher network. With 10,144 and 7,766 parameters for PeMSD7 and PeMSD8 ( Table 1 and Figure 1 ), it maintains accuracy while reducing complexity and computational requirements. This lightweight solution is efficient for traffic prediction, especially in resource-constrained scenarios.
Figure 1: The illustration features both our student and teacher models, demonstrating the application of our cost functions to the spatio-temporal graph convolutional network (ST-GCN)
3.2 Knowledge Distillation
We utilized two knowledge distillation techniques. The initial approach, known as response-based distillation, involves transmitting crucial information from the teacher network’s outputs to the student network. This leads to the development of a simpler and faster model while simultaneously preserving performance. The second method, feature-based knowledge distillation, focuses on hidden layers’ knowledge, encompassing spatial and temporal correlations among graph nodes (Figure 1). Knowledge distillation proves valuable for transferring knowledge from complex networks to simpler ones, particularly in resource-constrained environments requiring rapid algorithm execution.
@@ 3.2.1 Response-based Distillation
Leveraging the L2 and Kullback-Leibler (KL) divergence metrics, these functions measure the differences between the teacher and student network outputs [27, 28, 30, 29, 25]. The student network endeavors to enhance its accuracy compared to the teacher by minimizing these error functions. Equations (2) and (3) employ the L2 error function to quantify the difference between the outputs of the teacher and student networks for each node. This function assesses accuracy by summing the squared differences between the values of each network. The KL divergence error function in Equation 3 gauges the disparity between the probability distributions of the outputs of the teacher and student networks.
@@
In these equations, $||\cdot||_{2}$ denotes the L2 norm. The variables $N$, $y^{s}$, and $y^{t}$ represent the number of nodes, outputs of the student network, and outputs of the teacher network, respectively. KL is the Kullback-Leibler divergence metric (Equation 1), $T_{bi}$ is the target data value for batch $b$, and $\beta$ is an adjustable coefficient.
These cost functions consider both the teacher and target data simultaneously. However, we propose an alternative approach described by Equations (4) to (6) to decide when to utilize teacher predictions and when to use target data during student training. This decision is based on evaluating the difference between teacher predictions and target data. If this difference exceeds a threshold ($\alpha_{1}$), indicating potential inaccuracy, we incorporate the teacher’s prediction during student training for that node. Otherwise, we use the target data itself. This decision is made to account for potential noise in the target data, making the teacher’s prediction more valuable in such cases.
@@
In these equations, $D^{t}_{bi}$ represents the absolute differences between teacher predictions and target data for each node in batch $b$. After normalization (Equation 5), each element is compared with the threshold $\alpha_{1}$ to identify potentially noisy data. The loss function $L_{ORD}$ is then computed by summing these values for each node across all batches. You can observe the representation of this cost function in Figure 1.
3.2.2 Feature-based Knowledge Distillation in Hidden Layers
Knowledge distillation from hidden layers simplifies the training process for deep and complex neural networks, making it less intricate and time-consuming [31]. This method enables the training of simpler models that preserve meaningful information from the data and leverage the hidden knowledge of more complex models. Particularly beneficial in resource-constrained scenarios, this approach optimizes model training with limited resources. In the following subsections, we elaborate on our cost functions designed to capture both spatial and temporal correlations between the teacher and student networks. These cost functions aim to align the output of corresponding layers, fostering a close relationship in both spatial and temporal aspects. By addressing the intricacies of temporal correlation in temporal correlation distillation (described in equations (7) and (8)), and recognizing the resemblance between spatial and temporal layers in spatial correlation distillation (explained through equations (9) and (10)), our approach seeks to enhance the performance and knowledge transferability of the student network.
Temporal Correlation Distillation We introduce the cost function $L_{\text{TCD}}$ (as defined in equations (7) and (8) and illustrated in Figure 1 ) to ensure that the output of temporal layers in the student network closely aligns with the corresponding layers in the teacher network. The objective of $L_{\text{TCD}}$ is to enable the temporal layers in the student network, which have fewer parameters compared to equivalent layers in the teacher network, to perform similarly. By quantifying differences in temporal correlation values between nodes, this cost function guides the student network toward approximating these values in the teacher network.

### v4

@@ Abstract
Traffic prediction , Spatial-temporal graph knowledge distillation , Spatial-temporal graph neural network pruning , Model compression, Teacher-student architecture
@@ 1 Introduction
Traffic prediction is a critical task in urban planning, logistics, and transportation management Yu et al. [2018], Tedjopurnomo et al. [2022]. The data necessary for accurate prediction is typically sourced from monitoring stations strategically positioned throughout urban areas. These stations continuously collect a plethora of traffic-related parameters, including vehicle speed, traffic density, and flow rates, at regular intervals. This data is then structured into temporal graphs, with each node representing the traffic conditions observed at a specific monitoring station, and the edges indicating the connections between these stations Yu et al. [2018], Jiang and Luo [2022]. To leverage this wealth of data for predictive purposes, spatio-temporal graph neural networks (ST-GNNs) have emerged as a promising approach. These networks have the capability to analyze the temporal evolution of traffic patterns and make accurate predictions regarding future traffic conditions for each monitoring station within desired time intervals. However, a significant challenge arises when dealing with a large number of nodes, as it can lead to computationally expensive operations and demand substantial hardware resources Yu et al. [2018].
In response to this challenge, our research focuses on enhancing the execution time of ST-GNNs through the application of an integrated algorithm that combines knowledge distillation and network pruning Hinton et al. [2015], Zhang et al. [2019], Meng et al. [2019], Passban et al. [2021], Wang et al. [2020], Li et al. [2024], Ji et al. [2022], Li et al. [2022], Jiang and Luo [2022], Molchanov et al. [2019].
Previous works have not implemented an integrated algorithm to address the structure search for student networks by considering the capability of learning by filters using knowledge distillation. These studies typically consider the student structure static, or those that do address structure search for finding student structures do not mention using knowledge distillation to address the capability of filters to learn from the teacher. Instead, they simply use task-specific cost functions. In contrast, we address both aspects within a single algorithm, integrating them seamlessly in one training phase. Our approach uses knowledge distillation to train a simplified ”student” network with fewer parameters, helping it to understand complex spatial and temporal data by learning from a more complex ”teacher” network. To refine the student network’s architecture, we introduce a pruning algorithm that scores the capability of student filters in learning this information. Insignificant parameters are eliminated, and after each pruning stage, we fine-tune the network using knowledge distillation. This iterative process enhances the student network’s efficiency and learning capability, leading to improvements in execution time without compromising accuracy Hinton et al. [2015], Gou et al. [2021], Cheng et al. [2020], Yim et al. [2017], Yang et al. [2021]. By integrating pruning and knowledge distillation, our method effectively addresses the dual challenge of structuring student networks and enhancing their ability to learn complex data from the teacher model, resulting in a streamlined and effective training process.
The evaluation of our proposed approach is conducted using real-world traffic datasets, PeMSD7 and PeMSD8. Our results demonstrate that the student network trained using our cost function outperforms previous approaches in terms of both knowledge distillation effectiveness and execution time reduction. By jointly applying knowledge distillation and pruning, we assess the competency of network parameters, ensuring efficient model compression without sacrificing predictive accuracy.
@@ 2.1 Traffic Prediction and Spatio-Temporal Graph Neural Networks
Fast and accurate urban traffic prediction is vital for traffic control and management, especially for medium and long-term forecasts. Traditional methods often struggle with complexities in traffic flow and neglect spatial and temporal dependencies. To address this, deep networks capable of handling such data have been proposed.
In Ramakrishnan and Soni [2018], the significance of traffic network prediction and its applications, such as network monitoring, is discussed. Various recurrent neural network architectures, including standard RNN, LSTM, and GRU, are explored for traffic prediction. However, these models suffer from high execution times and difficulties in capturing long-range dependencies due to numerous parameters.
Bidirectional LSTM models are found to outperform unidirectional LSTM models in Abduljabbar et al. [2021], offering improved accuracy. Despite this, the high computational cost remains a significant drawback, making these models less suitable for real-time applications.
Spatio-temporal graph neural networks (ST-GNNs) are introduced to address these limitations by utilizing graph convolution layers for spatial processing and convolution layers for capturing temporal correlations. This approach offers reduced execution time and comparable or better accuracy.
In Guo et al. [2019], the attentional spatio-temporal graph convolutional network (ASTGCN) model effectively predicts traffic flow by leveraging spatial-temporal attention mechanisms and convolution layers. However, the model complexity and training time can still be high.
The spatio-temporal graph convolutional networks (STGCN) framework presented in Yu et al. [2018] shows enhanced training efficiency and captures comprehensive spatial-temporal correlations, outperforming other methods on real traffic datasets. Yet, it still faces challenges in scaling to very large networks and handling diverse urban environments.
Recent surveys such as Jiang and Luo [2022] and Jiang et al. [2023] provide comprehensive overviews of advancements in this field, highlighting the progress and persistent challenges such as handling heterogeneous data sources and improving model interpretability.
@@ 2.2 Knowledge Distillation
Knowledge Distillation, or Teacher-Student Learning, introduced in Hinton et al. [2015] in 2015, transfers knowledge from a complex ”teacher” network to a smaller ”student” network. This alignment of behavior and predictions improves student performance and reduces computational resources.
Further research extends this to graph and spatio-temporal networks. In Yang et al. [2020], a novel approach for knowledge transfer from a GCN teacher model to a GCN student model is introduced, leading to a smaller model with improved execution time, particularly suited for dynamic graph models. However, this method may struggle with the heterogeneity of traffic data.
In Jing et al. [2021], the objective is to train a simpler network through knowledge distillation from multiple GCN teacher models with different tasks, extending to distill spatial graph information from hidden convolutional layers. This approach empowers the student model to excel in various tasks without additional labeled data, but the integration of heterogeneous knowledge remains challenging.
In Bian et al. [2021], knowledge distillation is extended to ST-GNNs for modeling spatial and temporal data in human body position videos, utilizing various techniques, including minimizing loss, leveraging spatial-temporal relations, and using the gradient rejuvenation method to optimize the student model. Despite its success, adapting this approach to the diverse and dynamic nature of urban traffic data poses significant challenges.
@@ 2.3 Pruning and Fine-Tuning
In 1990, neural network pruning was first introduced by LeCun et al. [1989], offering a method to identify connections safe for pruning using the Hessian matrix. This reduces parameters in large networks, mitigating overfitting and resource requirements, crucial for deployment on constrained devices. Pruning has since become essential for compressing networks across diverse applications.
In 2019, Molchanov et al. [2019] proposed a method to estimate layer and weight importance, facilitating the removal of unimportant weights and network size reduction. Iterative pruning and retraining stages improve performance and reduce model complexity. However, maintaining performance while significantly reducing model size remains a persistent challenge.
@@ 3 Proposed Methods
Traffic prediction, involves predicting future values of key traffic conditions (e.g., speed or traffic flow) for the next $h$ time steps based on $M$ prior traffic observations. In this paper, we model the traffic network as graphs defined across time steps. At time step $t$, data from $N$ monitoring stations is represented as a graph $G_{t}=(V_{t},E,W)$, where $V_{t}\in\mathbb{R}^{N}$ contains features from $N$ nodes (e.g., road segments) with speed as the chosen criterion. The edge set $E$ depicts connections between stations, defined based on distance criteria, reflecting the influence of stations on each other. Weights $W\in\mathbb{R}^{N\times N}$ in the weighted adjacency matrix signify the relationships between stations, often determined by spatial disparities.
The ST-GCNYu et al. [2018], designed for traffic prediction, processes spatial and temporal data. Graph convolution layers capture spatial relationships and node features over time. Convolutional layers model temporal correlations, focusing on sequential patterns and changes between graphs. The architecture comprises two blocks with temporal and spatial layers, and an output block transforms information into the final output. The teacher and student networks share an identical number of blocks and layers. Any variations in parameters result from differences in channel numbers within the temporal and spatial layers of the hidden layer. The complex ST-GCN teacher network achieves high accuracy with 333,604 and 296,426 parameters for PeMSD7 and PeMSD8 datasets ( Table 1 and Figure 1 ). Despite accuracy, its high parameter count results in significant computational cost ( see Table 1, first row ). A proposed lighter model, the student network ( see Table 1, third row ) is a scaled-down version of the teacher network. With 10,144 and 7,766 parameters for PeMSD7 and PeMSD8, it maintains accuracy while reducing complexity and computational requirements. This lightweight solution is efficient for traffic prediction, especially in resource-constrained scenarios.
3.1 Knowledge Distillation
In this paper, we use offline distillation, wherein the student distills information from the teacher network after the teacher network has been completely trainedAsif et al. [2020], Mirzadeh et al. [2020], Mobahi et al. [2020]. We utilized two knowledge distillation techniques. The initial approach, known as response-based distillation, involves transmitting crucial information from the teacher network’s outputs to the student network. This leads to the development of a simpler and faster model while simultaneously preserving performance. The second method, feature-based knowledge distillation, focuses on hidden layer‘s knowledge, encompassing spatial and temporal correlations among graph nodes (Figure 1).
Figure 1: The illustration features both our student and teacher models, demonstrating the application of our cost functions to the spatio-temporal graph convolutional network (ST-GxCN)
@@ 3.1 Knowledge Distillation
PeMSD8 Models Parameters Hidden Blocks Channel Test Time (s) FLOPS Teacher 296,426 [1, 32, 64][64, 32, 128] 2.556 40,636,466,453 Pruning Base Model 39,290 [1, 8, 16][16, 8, 32] 0.810 5,659,617,749 Student 7,766 [1, 2, 4][4, 2, 8] 0.441 1,003,700,933
@@ 3.1.1 Response-based Distillation
Leveraging the L2 and Kullback-Leibler (KL) divergence metrics, these functions measure the differences between the teacher and student network outputs. The student network endeavors to enhance its accuracy compared to the teacher by minimizing these error functions. Equations (2) and (3) employ the L2 and KL error functions to quantify the difference between the outputs of the teacher and student networks for each node.
@@
In these equations, $||\cdot||_{2}$ denotes the L2 norm. The variables $N$, $y^{s}$, and $y^{t}$ represent the number of nodes, outputs of the student network, and outputs of the teacher network, respectively. KL is the Kullback-Leibler divergence metric (Equation (1)), $T_{bi}$ is the target data value for batch $b$, and $\beta$ is an adjustable coefficient. These cost functions consider both the teacher and target data jointly.
However, we propose an alternative approach described by Equations (4) to (6) to determine when to use teacher predictions versus target data during student training. This decision hinges on the difference between teacher predictions and target data. If this difference exceeds a threshold ($\alpha_{1}$), it indicates that even the more complex teacher model struggles to make an accurate prediction. In such cases, the simpler student model, with fewer parameters, is likely to face similar difficulties. Therefore, it is beneficial for the student to follow the teacher’s prediction, which tends to be smoother and less affected by noise. Otherwise, we use the target data directly. This method helps mitigate noise in the target data while leveraging the teacher’s more refined predictions to guide the student’s learning process effectively.
@@
In these equations, $D^{t}_{bi}$ represents the absolute differences between teacher predictions and target data for each node in batch $b$. After normalization (Equation (5)), each element is compared with the threshold $\alpha_{1}$ to identify potentially noisy data. The loss function $L_{ORD}$ is then computed by summing these values for each node across all batches. You can observe the representation of this cost function in Figure 1.
3.1.2 Feature-based Knowledge Distillation in the Hidden Layers
Knowledge distillation from hidden layers simplifies the training process for deep and complex neural networks, making it less intricate and time-consuming. This method enables the training of simpler models that preserve meaningful information from the data and leverage the hidden knowledge of more complex models. In the following subsections, we elaborate on our cost functions designed to capture both spatial and temporal correlations between the teacher and student networks. These cost functions aim to align the output of corresponding layers, fostering a close relationship in both spatial and temporal aspects.
Temporal Correlation Distillation We introduce the cost function $L_{\text{TCD}}$ (as defined in equations (7) and (8) and illustrated in Figure 1 ) to ensure that the output of temporal layers in the student network closely aligns with the corresponding layers in the teacher network. The objective of $L_{\text{TCD}}$ is to enable the temporal layers in the student network, which have fewer parameters compared to equivalent layers in the teacher network, to perform similarly.